OpenStudy (anonymous):

solve 2secx+tanx=3 on the interval xE[0,2pi]

OpenStudy (anonymous):

this is a pain.

OpenStudy (anonymous):

we have \[\frac{2}{\cos(x)}+\frac{\sin(x)}{\cos(x)}=\frac{2+\sin(x)}{\cos(x)}=3\] so \[2+\sin(x)=3\cos(x)\] \[3\cos(x)-\sin(x)=2\] \

OpenStudy (anonymous):

now we have to write \[3\cos(x)-\sin(x)\] as a single function of sine, we get something like \[\sqrt{10}\sin(x+\tan^{-1}(3))\]

OpenStudy (anonymous):

no that is a mistake

OpenStudy (anonymous):

\[-\sqrt{10}\sin(x-\tan^{-1}(3))=2\] i think it is right now

OpenStudy (anonymous):

i wonder if there is an easier way, but i don't see it. you can divide by \[-\sqrt{10}\]and take the inverse sine of the result

OpenStudy (anonymous):

let me try something else

OpenStudy (anonymous):

no just going around in circles. the above will do it. do you have another method you are supposed to use?